Obtain P(A | B) for the indicated probability. Round your answer to 3 decimal places. P(A) = 0.5, P(AC) = 0.5, P(B | A) = 0.3, P(B | AC) = 0.5.
Answer #1

P(B and A) = P(B | A) * P(A) = 0.3 * 0.5 = 0.15

P(B and Ac) = P(B | Ac) * P(Ac) = 0.5 * 0.5 = 0.25

P(B) = P(B and A) + P(B and Ac) = 0.15 + 0.25 = 0.4

P(A | B) = P(A and B) / P(B) = 0.15 / 0.4 = 0.375 (ans)